# Learning Keypoints for Robotic Cloth Manipulation using Synthetic Data

### v1

@@ Abstract
Assistive robots should be able to wash, fold or iron clothes. However, due to the variety, deformability and self-occlusions of clothes, creating general-purpose robot systems for cloth manipulation is challenging. Synthetic data is a promising direction to improve generalization, though its usability is often limited by the sim-to-real gap. To advance the use of synthetic data for cloth manipulation and to enable tasks such as robotic folding, we present a synthetic data pipeline to train keypoint detectors for almost flattened cloth items. To test its performance, we have also collected a real-world dataset.
We train detectors for both T-shirts, towels and shorts and obtain an average precision of 64.3%. Fine-tuning on real-world data improves performance to 74.2%. Additional insight is provided by discussing various failure modes of the keypoint detectors and by comparing different approaches to obtain cloth meshes and materials. We also quantify the remaining sim-to-real gap and argue that further improvements to the fidelity of cloth assets will be required to further reduce this gap. The code, dataset and trained models are available at https://github.com/tlpss/synthetic-cloth-data.
@@ I Introduction
In this paper, we want to enable generic robot folding. To this end, we generate synthetic data to learn keypoints on various cloth items. We focus on detecting non-occluded keypoints from an RGB image of almost flattened clothes for various cloth categories. Using these keypoints, tasks such as folding flattened clothes can be achieved with scripted motions [12, 13, 5]. Several prior works use synthetic data for cloth manipulation [14, 15, 5] but to the best of our knowledge, none have tackled this specific setting. Furthermore, we are also the first to compare the performance of different procedures to obtain cloth meshes and materials.
@@
We generate synthetic data for three different cloth categories: T-shirts, towels and shorts. This process has three stages: We first procedurally create cloth meshes. In the second stage, these meshes are dropped and deformed to mimic the output of current unfolding systems, for which we use Nvidia Flex [16]. Finally, we render images of the clothes and generate the corresponding annotations. To provide real-world training data and to thoroughly evaluate the performance of the keypoint detectors, we have also gathered a dataset of almost 2,000 images containing more than 100 cloth items.
We find that keypoint detectors trained on our synthetic data have a mean average precision of 64.3 % on our test dataset, which is an improvement over the baseline trained only on real data. Fine-tuning using real data further improves performance to 74.2%. These results show that our synthetic data pipeline can be used to build state estimators for robotic cloth folding systems. Additionally, we compare several different procedures to generate cloth meshes and materials where we find that the best results are obtained by using random materials and single-layer cloth meshes, even though this decreases fidelity. Finally, we quantify the remaining sim-to-real gap and argue that this gap can only be closed by further advancements in cloth asset generation and simulation.
@@ II-A Robotic Cloth Manipulation
Given flattened clothes, tasks such as folding can be tackled using an appropriate representation and a scripted policy [5, 13]. Several methods have been explored to create such state representations for clothes, including template fitting [12, 4], edge detection [18] and semantic keypoint detection [19, 5, 14, 20]. Many of these works use depth images [14, 18, 20], as depth images are in theory invariant to cloth materials. However, they also discard useful information (such as seams) and additionally can lack the required precision [18].
In this work, we focus on learning keypoints on RGB images of almost-flattened clothes that are lying on a surface. By doing so, we aim to facilitate the creation of cloth manipulation systems that operate on the output of an unfolding system and can generalize to arbitrary clothes and environments.
@@ II-B Synthetic data for Robotic Cloth Manipulation
To train and evaluate the keypoint detectors, we have created a dataset of (almost) flattened clothes. We aim to enable folding clothes that have been previously unfolded (flattened) and because current unfolding pipelines do not yet produce perfectly flattened results [5, 6], we have applied small deformations to the clothes. We refer to this dataset as the aRTF Clothes dataset, for almost-ready-to-fold clothes. There have already been some efforts to collect datasets for robotic cloth manipulation [4, 26, 27], but to the best of our knowledge, this dataset is the first to provide labeled images of almost-flattened cloth items in household settings.
@@ III-A Dataset Capturing
To collect the images we listed 14 household scenes containing a flat surface that can be used to fold clothes. We then collected several clothing items by asking lab members to bring them. For the towels, we included some pieces from the household cloth dataset [28].
For each image, we started by laying the cloth item on the folding surface. Based on examples from state-of-the-art unfolding pipelines [4, 5, 6], we distilled a number of failure modes and linked these to deformations that we randomly applied to the flat cloth items. These deformations include a random pinch to create wrinkles, or folding a side or corner of the cloth upwards or downwards.
We collected the images using a ZED2i RGB-D camera. For a number of scenes we used a smartphone alongside the ZED camera, to incorporate camera and lens diversity. The camera was positioned to mimic an ego-centric perspective. The distance to the folding surface is between 0.5 and 1 m.
The dataset contains a total of 1,896 images and is summarized in Table I. Collecting these images took 2 persons about 15 hours. Some example images from the dataset can be found in Fig. 1.
@@ III-B Annotating
For each cloth category, we defined a number of semantic locations (keypoints) and annotated these on each image. We mostly use the same semantic locations as Miller et al. [29]. The semantic locations can be seen in the examples in Fig. 1.
We also labeled the segmentation mask and bounding boxes, for which we use Segment Anything (SAM) [30]. We have found SAM regularly fails with a single point, but produces satisfactory masks when provided with or 2 to 3 points on the cloth surface. In total, this labeling effort took a single person about 3 days.
@@ IV Procedural Data Generation Pipeline
In this section, we describe the synthetic data pipeline we have created to generate synthetic images of clothes. In general, generating synthetic data entails two steps: obtaining the assets (meshes, materials…) and specifying how these assets should be used to create 3D scenes, from which the desired images and their corresponding annotations can then be generated. Powerful tools such as Blender[23] have enabled the latter step. The main challenge typically lies in gathering or creating the required assets, and this typically comes with a trade-off between realism and diversity. As we are dealing with deformable objects, generating the assets is even more complicated: we need not only to obtain cloth meshes but also to generate the desired distribution of configurations for them, i.e., we need to deform them. The next sections will discuss in more detail how we generate cloth meshes and materials, deform the clothes using Nvidia Flex [16], and use Blender [23] to generate images and the corresponding annotations.
@@ IV-A1 Cloth meshes
The first step is to obtain flat cloth meshes. We generate a series of 2D boundary vertices using a template for each cloth type. These templates are inspired on [29]. We then connect the vertices using bezier curves to create single-layer meshes. These bezier curves are used to create the neck of a T-shirt and to better mimic real cloth items by slightly curving edges. For the same reason, we also round the corners of the single-layer mesh. The parameters of the skeleton, the bezier curves and the rounding radii are sampled from manually tuned ranges. Finally, the meshes are triangulated such that they have edge lengths of at most 1cm and UV maps are generated. We keep track of the vertices that correspond to each semantic location to automatically label the keypoints later on.
@@ IV-A2 Cloth Materials
Faithfully reproducing cloth materials requires both mimicking fabrics and producing realistic color patterns. At the same time, we also need these materials to provide sufficient diversity to enable the generalization of the models trained on the data. This combination tends to be infeasible. We have experimented with various procedures (see Section VI-B) to generate the cloth materials and have found that using random textures from PolyHaven [31] provides the best results. To increase diversity, we mix these textures with a random color.
@@ IV-B Deforming the Meshes
mAP2,4,8 (normal-$\uparrow$) mAP2 (normal-$\uparrow$)
@@
T-shirt 54.4 58.2 69.1 30.5 43.8 51.9
Towel 74.4 83.2 88.6 65.9 77.7 83.5
Shorts 50.7 51.4 64.9 37.5 44.0 55.7
All 59.8 64.3 74.2 44.6 55.2 63.7
@@ IV-C Scene Composition
We sample an environment texture from PolyHaven [31] to create complex scene lighting. At the same time, we use this texture to create a scene background. Thereafter, a rectangular surface is added to the scene to mimic the folding surface. We sample a material from PolyHaven [31] for this plane and randomize its base color to increase diversity and avoid a bias towards certain colors. We then put a cloth mesh on top of the table and solidify the mesh to add some thickness to the layers. Afterwards, we sample a cloth material and apply it to the mesh. To make the keypoint detectors more robust, we place a number of distractor objects on the surface. We sample the objects from the Google Scanned Objects dataset [33]. Next, we randomize the camera position in a spherical cap around the cloth center and point the camera towards the cloth. At this point we render a single image using Cycles, Blender’s Physically Based Renderer, and create a bounding box, segmentation mask and 2D keypoint annotations. For the keypoints, we use raycasting to determine if the vertex that corresponds to the keypoint is occluded. Inspired by [10], we consider a keypoint as visible if any vertex in its N-ring neighborhood is visible. This corresponds better to the human labels on the aRTF dataset.
@@ V Keypoint Detection
This section provides more details about the keypoint detectors. We describe the model used to detect keypoints and the average precision metrics we use to quantify the performance of the keypoint detectors. We also describe how we ordered the keypoints in image space to deal with the symmetries of cloth objects.
@@ V-A Model and Training
We formulate keypoint detection as regression on 2D heatmaps as in [34]. With this formulation, keypoint coordinates are converted into Gaussian blobs on a 2D heatmap, and these heatmaps are then used target for pixel-wise regression. The keypoint model uses a Unet [35] inspired architecture. The encoder of the Unet is a pretrained MaxViT [36] nano model, obtained from the timm library [37].
As is common  [10, 8], we employ a set of image augmentations during training: random cropping, random Gaussian blurring and noise, and random color, contrast and brightness augmentations. Note that some of these effects could be added to the synthetic data pipeline, but we follow [10] by adding them as augmentations to increase flexibility.
@@
V-B Evaluation Metric
In this work, we want to enable folding. Hence, we only detect keypoints that are not occluded, as a robot can only interact with those parts of the cloth. This implies a varying number of detections for each keypoint category and hence we cannot directly use adistance-based metric such as the average pixel distance [14, 20]. Instead, we use the Average Precision (AP) to quantify the performance of the keypoint detectors. To distinguish between false positive and false negative keypoint predictions we use the L2 pixel distance [^1] COCO uses a non-linear transform of the L2 pixel distance: the object-similarity score (OKS). We directly use the L2 distance, since it relates more directly to real-world distances, which are typically the metric of interest for robotic manipulation..
As thresholds for the L2 distance, we use 2, 4 and 8 pixels. For the ZED 2i with an image resolution of 512x256 and given that the maximal distance between a cloth and the camera in the aRTF dataset is 1 meter (see Section I), this corresponds to real-world discrepancies between real and predicted keypoints of at most 1, 2 and 4 cm. The thresholds were chosen as follows: 1cm and 2cm are relevant for grasping as they are the half-width of regular fingertips and of custom fingertips used in [19]. 4cm is relevant for tasks such as ironing and serves as upper bound to distinguish between lack of precision and lack of semantic knowledge for the keypoint detectors. Similar in spirit to the COCO metrics [38], we report the mean average precision over all thresholds (mAP2,4,8 ) and the AP on the most strict threshold (mAP2 ).
@@ V-C Dealing with symmetry
Clothes have (apparent) symmetries that need to be handled carefully when detecting keypoints, as the model cannot be expected to distinguish between a motion of the camera versus a motion of the cloth object. Therefore we need to handle all symmetries in the Special Euclidean group SE(3) (which describes all rigid motions and hence all possible camera movements). One possible solution is to not distinguish between symmetric keypoints by labeling them identically [19], but this results in some loss of information: we can no longer infer the connectivity of the keypoints, which is useful if you want to e.g. grasp the two corners of an edge. Other solutions include using a loss function that is invariant across all symmetries [14], or selecting an arbitrary ordering in the image [5, 24]. We have taken the latter approach in this work and have adapted the labels of the synthetic data and the aRTF dataset as follows:
@@
T-shirts or shorts have no symmetries in SE3 (which does not include reflections). However, detecting the front/back side can be very ambiguous, even on real data, and therefore we treat this as a (rotational) symmetry. For T-shirts we take as left side, the side of which the waist keypoint is closest to the bottom left corner of the T-shirt’s bounding box. For shorts, we use the top left corner of the bounding box.
@@ VI Experiments
In this section, we train several keypoint detectors on different data sources. We evaluate the performance of all models on the test split of the aRTF dataset (see Table I) using the mean Average Precision (mAP) as metric. In Section VI-A we show that training on the synthetic data for three cloth categories results in better performance than training on the aRTF train split, and how fine-tuning on the real data improves performance significantly. The cloth mesh procedure and cloth materials described in Section 2 are compared against other alternatives in Sections VI-B and VI-C.

### v2

@@ Abstract
Assistive robots should be able to wash, fold or iron clothes. However, due to the variety, deformability and self-occlusions of clothes, creating robot systems for cloth manipulation is challenging. Synthetic data is a promising direction to improve generalization, but the sim-to-real gap limits its effectiveness. To advance the use of synthetic data for cloth manipulation tasks such as robotic folding, we present a synthetic data pipeline to train keypoint detectors for almost-flattened cloth items. To evaluate its performance, we have also collected a real-world dataset.
We train detectors for both T-shirts, towels and shorts and obtain an average precision of 64% and an average keypoint distance of 18 pixels. Fine-tuning on real-world data improves performance to 74% mAP and an average distance of only 9 pixels. Furthermore, we describe failure modes of the keypoint detectors and compare different approaches to obtain cloth meshes and materials. We also quantify the remaining sim-to-real gap and argue that further improvements to the fidelity of cloth assets will be required to further reduce this gap. The code, dataset and trained models are available here.
@@ I Introduction
In this paper, we want to enable generic robot folding. To this end, we generate synthetic data to learn keypoints on various cloth items after they haven been unfolded by a robot. Current unfolding systems are not perfect yet [3, 5, 6] and we therefore detect keypoints on both flattened and slightly deformed clothes. We refer to these cloth states as almost-flattened. We detect non-occluded keypoints on clothes of various categories from an RGB image. Using these keypoints, tasks such as folding flattened clothes can be achieved with scripted motions [12, 13, 5]. Several prior works use synthetic data for cloth manipulation [14, 15, 5] but to the best of our knowledge, none have tackled this specific setting. Furthermore, we are also the first to compare the performance of different procedures to obtain cloth meshes and materials.
@@
We generate synthetic data for T-shirts, towels and shorts. This process has three stages: We first procedurally create cloth meshes. In the second stage, these meshes are dropped and deformed to mimic the output of current robotic unfolding systems, using Nvidia Flex [16]. Finally, we render images of the clothes and generate the corresponding annotations. To provide real-world training data and to thoroughly evaluate the performance of the keypoint detectors, we have also gathered a dataset of almost 2,000 images.
We find that keypoint detectors trained on our synthetic data have a mean average precision of 64.3 % on our evaluation dataset, which is an improvement over the baseline trained only on real data. Fine-tuning using real data further improves performance to 74.2%. These results show that our synthetic data pipeline can be used to build state estimators for robotic cloth folding systems. Additionally, we compare several different procedures to generate cloth meshes and materials where we find that the best results are obtained by using random materials and single-layer cloth meshes, even though this decreases fidelity. Finally, we quantify the remaining sim-to-real gap and argue that this gap can only be closed by further advancements in cloth asset generation and simulation.
@@ II-A Robotic Cloth Manipulation
Given flattened clothes, tasks such as folding can be tackled using an appropriate representation and a scripted policy [5, 13]. Several methods have been explored to create such state representations for clothes, including template fitting [12, 4], edge detection [18] and semantic keypoint detection [19, 5, 14, 20]. Many of these works use depth images [14, 18, 20], as these only represent geometry and are not influenced by irrelevant environment elements such as lighting. However, depth images also discard useful information (such as color patterns, seams..) and can lack the required precision [18]. To avoid these limitations, we use RGB images in this work.
In this work, we detect keypoints of almost-flattened clothes that are lying on a surface. By doing so, we aim to facilitate the creation of cloth manipulation systems that operate on the output of an unfolding system and can generalize to arbitrary clothes and environments.
@@ II-B Synthetic data for Robotic Cloth Manipulation
To train and evaluate the keypoint detectors, we have created a dataset of almost-flattened clothes in which we mimic the output of current robotic unfolding systems [5, 6, 3] by applying small deformations to the clothes. We refer to this dataset as the aRTF Clothes dataset, for almost-ready-to-fold clothes. There have already been some efforts to collect datasets for robotic cloth manipulation [4, 26, 27], but to the best of our knowledge, this dataset is the first to provide annotated real-world of almost-flattened cloth items in household settings.
@@ III-A Dataset Capturing
We collected data in 14 distinct household scenes. The cloth items were provided by lab members, though we included towels from the household cloth dataset [28].
For each image, we started by laying the cloth item on the folding surface and then randomly applied some deformations. The deformations include a random pinch to create wrinkles, or folding a side or corner of the cloth upwards or downwards and were determined by listing failure modes of state-of-the-art unfolding pipelines [4, 5, 6].
We collected the images using a ZED2i RGB-D camera and a smartphone. The camera was positioned to mimic an ego-centric perspective at a distance between 0.5 and 1 m.
The dataset contains a total of 1,896 images and is summarized in Table I. Collecting these images took 2 persons about 15 hours. Fig. 1 shows some examples from the dataset.
@@ III-B Annotating
For each cloth category, we defined a number of semantic locations (keypoints) and annotated these on each image. We use the same semantic locations as Miller et al. [29], except for a single semantic location in the neck opening of T-shirts that we have not used in our work. The semantic locations can be seen in the examples in Fig. 1.
We also labeled the segmentation mask and bounding boxes and use Segment Anything (SAM) [30] for pre-labeling the masks. In total, this labeling effort took a single person about 3 days.
@@ IV Procedural Data Generation Pipeline
In this section, we describe the synthetic data pipeline we have created to generate synthetic images of clothes. In general, generating synthetic data entails two steps: obtaining the assets (meshes, materials…) and specifying how these assets should be used to create 3D scenes, from which the desired images and their corresponding annotations can then be generated. Tools such as Blender[23] have enabled the latter step. The main challenge typically lies in gathering or creating the required assets, and this typically comes with a trade-off between realism and diversity. As we are dealing with deformable objects, generating the assets is even more complicated: we need not only to obtain cloth meshes but also to generate the desired distribution of configurations for them, i.e., we need to deform them. The next sections will discuss in more detail how we generate cloth meshes and materials, deform the clothes using Nvidia Flex [16], and use Blender [23] to generate images and the corresponding annotations.
@@ IV-A1 Cloth meshes
The first step is to obtain flat cloth meshes. We generate a series of 2D boundary vertices using a template for each cloth type. These templates are inspired on [29]. We then connect the vertices using bezier curves to create single-layer meshes. These bezier curves are used to create the neck of a T-shirt and to better mimic real cloth items by slightly curving edges. For the same reason, we also round all corners of the single-layer mesh. The parameters of the skeleton, the bezier curves and the rounding radii for all corners are sampled from manually tuned ranges. Finally, the meshes are triangulated such that they have edge lengths of at most 1cm and UV maps are generated. We keep track of the vertices that correspond to each semantic location to automatically label the keypoints later on.
@@ IV-A2 Cloth Materials
Faithfully reproducing cloth materials for rendering requires both mimicking fabrics and producing realistic color patterns. At the same time, we also need these materials to provide sufficient diversity to enable the generalization of the models trained on the data. This combination tends to be infeasible. We have experimented with various procedures (see Section VI-B) to generate the cloth materials and have found that using random textures from PolyHaven [31] provides the best results. To increase diversity, we mix these textures with a random color.
@@ IV-B Deforming the Meshes
mAP2,4,8 ($\uparrow$) AKD($\downarrow$)
@@
T-shirt 54.4 58.2 69.1 14.2 14.0 8.3
Towel 74.4 83.2 88.6 12.5 13.1 6.8
Shorts 50.7 51.4 64.9 23.6 27.3 11.2
All 59.8 64.3 74.2 16.8 18.1 8.7
@@ IV-C Scene Composition
We sample an environment texture from PolyHaven [31] to create complex scene lighting. At the same time, we use this texture to create a scene background. Thereafter, a rectangular surface is added to the scene to mimic the folding surface. We sample a material from PolyHaven [31] for this plane and randomize its base color to increase diversity and avoid a bias towards certain colors. We then put a cloth mesh on top of the table and solidify the mesh to add some thickness to the layers. Afterwards, we sample a cloth material and apply it to the mesh. To make the keypoint detectors more robust, we place a number of distractor objects on the surface. We sample the objects from the Google Scanned Objects dataset [33]. Next, we randomize the camera position in a spherical cap around the cloth center and point the camera towards the cloth. At this point we render a single image using Cycles, Blender’s physically based renderer, and create a bounding box, segmentation mask and 2D keypoint annotations. For the keypoints, we use raycasting to determine if the vertex that corresponds to the keypoint is occluded. Inspired by [10] we have tried to make the synthetic labels resemble the human labels in the aRTF dataset, which are not necessarily similar to the ground truth labels. We consider a keypoint as visible if any vertex in its 2-ring neighborhood is visible in the image, as we empirically observed this to correspond better to our human annotations.
@@ V Keypoint Detection
This section provides more details about the keypoint detectors. We describe the model used to detect keypoints and the metrics we use to quantify the performance of the keypoint detectors. We also describe how we ordered the keypoints in image space to deal with the symmetries of cloth objects.
@@ V-A Model and Training
We detect the keypoints on the full-resolution images and formulate keypoint detection as regression on 2D heatmaps as in [34]. We convert keypoint coordinates into Gaussian blobs on a 2D probability heatmap, and then use these heatmaps (one for each keypoint category) as target for pixel-wise logistic regression using a binary cross entropy loss. The keypoint model uses a Unet [35] inspired architecture and has N full-resolution heatmaps as output. The encoder of the Unet is a pretrained MaxViT [36] nano model, obtained from the timm library [37]. Keypoints are extracted from the predicted heatmaps by searching for local maxima in a 3x3 pixel grid with a threshold of 0.01 on the probability. Downstream applications can then select the keypoint with the highest probability or use the multi-modal heatmaps to reason more globally over the predictions.
As is common  [10, 8], we employ a set of image augmentations during training: random cropping, random Gaussian blurring and noise, and random color, contrast and brightness augmentations.
@@
V-B Evaluation Metrics
Mean Average Precision. In this work we want to enable folding and hence we only detect keypoints that are not occluded, as a robot can only interact with those parts of the cloth. This implies a varying number of detections for each keypoint category. Furthermore our model can predict an arbitrary amount of keypoints, as discussed in Section V. Therefore, we use the Average Precision (AP) as in COCO [38] to measure the performance of our models. For each heatmap channel, we classify all predictions as false or true positives using the L2 distance to the ground truth keypoints with thresholds of 2, 4 and 8 pixels. For the aRTF dataset and our image resolution of 512 by 256 pixels, this corresponds to real-world distances between real and predicted keypoints of at most 1, 2 and 4 cm respectively. We report the mean average precision over all thresholds (mAP2,4,8 ). Note that unlike in COCO’s object-keypoint similarity (OKS), we do not scale the distance according to the object size as we believe that absolute errors are more appropriate for robotic manipulation: it does not matter if a T-shirt is a large or small, if you want to grasp it at a specific location you need the same accuracy for both.
Average keypoint Distance. Next to this detection measure, we also employ a distance-based metric to quantify performance as in [20, 14]. More specifically, we report the average keypoint distance (AKS), measured in pixels, between the predicted (if any) keypoint with the highest probability and the non-occluded ground truth keypoints.
Figure 3: Illustration of failure modes of the detectors. From left to right, the columns show ground truth keypoints and all predicted keypoints of detectors trained on real data, synthetic data and on both. The first row illustrates how the real baseline produces incomplete and inconsistent keypoints, whereas the second row shows how the detectors still struggle with folds. In the third row, the detectors are confused by an open zipper and mistake it for a leg of the shorts. The keypoint colors encode their category for each cloth type.
@@ V-C Dealing with symmetry
Clothes have (quasi-) symmetries, which means they can have the same appearance on an image for different configurations (e.g. rotating a white towel on a table by 180 degrees will result in the same image). These symmetries must be handled carefully because the keypoint detector cannot distinguish between them.
One possible solution is to not distinguish between symmetric keypoints by labeling them identically [19], but this results in some loss of information: we can no longer infer the connectivity of the keypoints, which is useful if you want to e.g. grasp the two corners of an edge. Other solutions include using a loss function that is invariant across all symmetries [14], or selecting an arbitrary ordering in the image [5, 24]. A thorough evaluation of these approaches would be out of scope and in this work we have applied an ordering to the keypoints to deal with symmetry. To do so, we have adapted the labels of the synthetic data and the aRTF dataset as follows:
@@
T-shirts or shorts have no rigid motions symmetries. However, detecting the front/back side can be very ambiguous, even on real data, and therefore we treat this as a rotational symmetry. For T-shirts we take as left side, the side of which the waist keypoint is closest to the bottom left corner of the T-shirt’s bounding box. For shorts, we use the top left corner of the bounding box.
@@ VI Experiments
In this section, we train several keypoint detectors on different data sources. We evaluate the performance of all models on the test split of the aRTF dataset (see Table I) using the mean Average Precision (mAP) as metric. In Section VI-A we show that training on the synthetic data (sim-to-real) for three cloth categories results in better performance than training on the aRTF train split (real-to-real), and how fine-tuning on the real data (sim+real-to-real) improves performance significantly. The cloth mesh procedure and cloth materials described in Section 2 are compared against other alternatives in Sections VI-B and VI-C.